OpenStudy (anonymous):

Find the x- and y-intercepts -4x-5y=30

OpenStudy (anonymous):

to find the x int set y to zero and solve for x. to find the y int set x to zero and solve for y

OpenStudy (anonymous):

x int: -4x-5(0)=30 -4x=30, divide both sides by -4 and you get: x=-15/2, so xint is (-15/2,0) y int: -4(0)-5y=30 -5y=30, divide both sides by -5 and you get y=-6, so y int is (0,-6)

OpenStudy (anonymous):

please click the good andswer button if my answers satify you